Grade 12Analytical Geometry

If the radius of the circle having maximum size passing through (2,4) and touching both the coordinate axes is

Ganesh
9 Years agoGrade 12

1 Answer

Himalaya Nirjhar Datta
9 Years ago
The question is a bit ambiguous but if we find the radius of the circle what is meant by its maximum size then the answer is 2 unit. Way to solve:Consider a line per perpendicular to x-axis which passes from the point (2,4).Then the point where it meets x-axis and the (2,4) point forms the diameter of the required circle.Radius = 1/2 distance between these two points = 1/2 × 4= 2 units.Please feel free to post any reply ( criticism or constructive) .